OpenStudy (anonymous):

You roll two dice. What is the probability of rolling a 3 or a 5 on the second die given that you rolled an odd number on the first die?

OpenStudy (anonymous):

How did you solve that? What formula did you use please?

OpenStudy (anonymous):

rolling a 3 or a 5 on the second die given that you rolled an odd number on the first die = {(1,3),(3,3),(5,3),(1,5),(3,5),(5,5)} ===> p ( the event) = 6/36 = 1/6

OpenStudy (anonymous):

all outcomes od rolling 2 dies = 36 , and the outcomes that meet the conditions =6 ===> p(the required event) =6/36 = 1/6

OpenStudy (anonymous):

how do you account for the double number combinations? I totally get lost with the probability stuff!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

ypu are welcom

OpenStudy (anonymous):

you are welcom again.